OpenStudy (anonymous):

i need help: find the derivative dy/dx. y= x^(1/2)*(e^(x^2))*(x^2+1)^10

OpenStudy (anonymous):

is it \[ y =\sqrt x e^{x^2} ( x^2+1)^{10} \]

OpenStudy (anonymous):

If it is use, the product rule for derivative \[g= u v w\\ g'= u' vw + v' uw + w' u v \]

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

So apply the product rule above and you will be done.

OpenStudy (anonymous):

but how to do with (x^2+1)^10

OpenStudy (anonymous):

20 x ( x^2 +1)^9

OpenStudy (anonymous):

Your final answer should look like \[ \frac{e^{x^2} \left(x^2+1\right)^{10}}{2 \sqrt{x}}+2 e^{x^2} x^{3/2} \left(x^2+1\right)^{10}+20 e^{x^2} x^{3/2} \left(x^2+1\right)^9 \]

OpenStudy (anonymous):

thanks i see

OpenStudy (blockcolder):

An alternative in case you encounter some more of these, plus quotients (probably): Take the ln of the function you're differentiating and use properties of ln to simplify: \[y =\sqrt x e^{x^2} ( x^2+1)^{10}\\ \ln{y}=\frac{1}{2}\ln{x}+x^2+10\ln(x^2+1)\] Differentiate wrt x and multiply by y: \[\frac{y'}{y}=\frac{1}{2x}+2x+\frac{20x}{x^2+1}\\ y'=\sqrt x e^{x^2} ( x^2+1)^{10}\left( \frac{1}{2x}+2x+\frac{20x}{x^2+1}\right)\]